# You May Use the Same Channel Knowledge Map for Environment-Aware NLoS Sensing and Communication

### v2

@@ I Introduction
Significant progress has been made for ISAC systems, particularly in line-of-sight (LoS) scenarios, where the sensing targets are directly visible to both the transmitter and receiver. Various ISAC techniques have been developed, like joint waveform design, beamforming optimization, and interference management to balance radar-like sensing with data transmission, achieving high localization accuracy and spectral efficiency in LoS conditions [13, 14, 15, 16, 17]. For instance, methods such as time-of-arrival (ToA) estimation and angle-based localization leverage direct LoS paths to estimate target locations [6, 18]. However, in complex environments such as urban low-altitude airspace, LoS paths are frequently blocked. As a result, sensing in such environments can only rely on pure non-line-of-sight (NLoS) signal paths, posing a much greater challenge to map the estimated sensing parameters to the location of the sensing targets as more unknowns are involved [19, 20]. Some techniques have been proposed to achieve NLoS localization, including fingerprinting [21], single-base-station (BS) localization exploiting near-field scatterers [22], and reconfigurable intelligent surface (RIS)-aided methods [23]. However, they cannot be applied for sensing since unlike localization tasks, the sensing targets are typically non-cooperative in the sense that they only passively reflect the signals rather than transmitting or receiving them. Specifically, fingerprinting based localization works by building a database that maps user equipment (UE) location to its wireless communication channel signature such as received signal strength (RSS). However, this approach is unsuitable for sensing. The fundamental reason is that a fingerprint database characterizes the direct channel between UE and the anchor nodes, whereas sensing relies on the echoes from the passive sensing target. Therefore, the fingerprint database built for localization cannot be used for wireless sensing. Instead, sensing would require its own “sensing database” that fingerprint target-specific echo characteristics. Creating such databases is quite challenging, if not impossible, as different types of targets (e.g., vehicles, pedestrians) would need a distinct and comprehensive map of its potential echo signatures. On the other hand, single-BS near-field localization studied in [22] first senses near-field scatterers to acquire environment priors, and then leverages the location of near-field scatterers to localize UEs without relying on LoS link. However, it relies on near-field assumptions, limiting applicability in far-field ISAC. In addition, RIS-aided methods create virtual LoS relationships for localization by deploying reflective surfaces, they require additional costly and extensive hardware deployments. In fact, rigorous theoretical studies have proven that, without environment or NLoS prior knowledge, NLoS paths cannot provide beneficial information for localization [24]. This demonstrates the imperative need to develop an innovative framework that can leverage environment prior knowledge to tackle the challenging problem of NLoS sensing for ISAC systems.
@@
The remainder of this paper is organized as follows. Section II presents the system model for the ISAC framework. Section III discusses the motivation for CKM-enabled sensing, while Sections IV details the mechanisms for the proposed CADM-enabled NLoS sensing, respectively. Section V provides simulation results, and Section VI concludes the paper.
@@ II System Model
where $c$ is the speed of light. As a result, after estimating the parameters $\hat{\theta}$ and $\hat{\tau}$ based on the echo signal $\mathbf{y}(t)$ in (1) with variety algorithms reviewed in [8], the location of the sensing target $\mathbf{x}$ can be directly obtained as
@@
which ignores a significant number of potential paths. In contrast, the proposed model in (4) accounts for all $L=L^{\prime 2}$ possible path combinations, the received signal in (4) can be written as
@@
To ultimately achieve the sensing goal, these parameters in $\mathbf{z}$ must be mapped to the location $\mathbf{x}$ of the sensing target. In LoS scenarios, this mapping can be directly achieved with the closed-form expression (3). However, such a mapping is no longer valid for the challenging scenario where LoS is absent, since NLoS paths dependent on the environment, denoted as $E$, as shown in Fig. 1. Therefore, when only NLoS paths are available, without any prior knowledge about the environment, determining the sensing target location $\mathbf{x}$ based on the sensed channel parameters $\mathbf{z}$ is an underdetermined problem. In fact, even in mixed scenarios with both LoS and NLoS paths, conventional methods typically employ techniques such as NLoS mitigation or path selection to isolate the LoS path, discarding NLoS paths that contain valuable target-related information [47]. This results in information loss, limiting the sensing accuracy. To utilize the target-related information in NLoS paths, which depends on $E$, an environment-aware sensing model that incorporates $E$ is necessary to leverage both LoS (if exists) and NLoS paths for sensing target localization.
@@
The problem (10) can be solved by classic optimization method such as gradient descent method. The update rule for the target location at the $(k+1)$th iteration is given by
@@ III Environment-aware ISAC via CKM
CKM was originally developed to infer prior channel knowledge based on UE locations for communication purposes. In this context, the sensing target location $\mathbf{x}$ was not involved. CKM offers a method to bypass the requirement for explicit modeling of $E$ or $f(\cdot)$, circumventing the complexities associated with environmental characterization and function derivation. Extending this framework to map sensing target locations $\mathbf{x}$ to channel knowledge $\mathbf{z}$ faces new challenge and opportunity, which motivates the exploration of CKM-enabled sensing in this work. In the following, we present the motivation behind this extension, detailing its core mechanism for enhancing environment-aware sensing capabilities.
Let $\mathbf{Q}=[\mathbf{q}_{1},\ldots,\mathbf{q}_{Q}]\in\mathbb{R}^{2\times Q}$ and $\mathbf{W}=[\mathbf{w}_{1},\ldots,\mathbf{w}_{Q}]\in\mathbb{R}^{D\times Q}$ denote the $Q$ historically visited UE locations and their corresponding observed $D$-dimension channel knowledge data, respectively [26]. These location-channel knowledge pairs $(\mathbf{Q},\mathbf{W})$ are utilized to learn prior channel knowledge, forming the basis of CKM. CKM can thus be expressed as
@@
Figure 3: A
n illustration of CKM-enabled ISAC that utilizes the same CKM $\mathcal{M}$ developed for communication to achieve environment-aware sensing as well. The CKM, by learning the “location-to-channel” mapping ($P_{\mathcal{M}}(\mathbf{w}|\mathbf{q})$) from data, serves as an implicit representation of the environment $E$. This allows the framework to “skip” the computationally expensive step of explicit physical environment reconstruction and directly provide the sensing likelihood $P(\hat{\mathbf{z}}|\mathbf{x})$ for sensing target localization.
@@ IV-A Relationship between sensing response and communication channel
$\mathbf{y}_{l_{R},l_{T}}(t)$ $=\mathbf{h}_{R,l_{R}}(t)*\big(\beta(\theta_{l_{T}},\theta_{l_{R}})y_{T,l_{T}}(t)\big)$ (27)
$=\big(\alpha_{l_{R}}\mathbf{a}(\theta_{l_{R}})\delta(t-\tau_{l_{R}})\big)*$
$\qquad\big(\beta(\theta_{l_{T}},\theta_{l_{R}})\alpha_{l_{T}}\mathbf{a}^{H}(\theta_{l_{T}})\mathbf{s}(t-\tau_{l_{T}})\big)$
$=\beta(\theta_{l_{T}},\theta_{l_{R}})\alpha_{l_{R}}\alpha_{l_{T}}\mathbf{a}(\theta_{l_{R}})\mathbf{a}^{H}(\theta_{l_{T}})$
$\qquad\times\mathbf{s}\big(t-(\tau_{l_{R}}+\tau_{l_{T}})\big).$
@@
$=\sum_{l_{R}=1}^{L^{\prime}}\sum_{l_{T}=1}^{L^{\prime}}\beta(\theta_{l_{T}},\theta_{l_{R}})\alpha_{l_{R}}\alpha_{l_{T}}\mathbf{a}(\theta_{l_{R}})\mathbf{a}^{H}(\theta_{l_{T}})$
$\qquad\times\mathbf{s}\big(t-(\tau_{l_{R}}+\tau_{l_{T}})\big)+\mathbf{n}(t)$
$=\mathbf{H}(t)*\mathbf{s}(t)+\mathbf{n}(t),$
@@
$\times\mathbf{a}^{H}(\theta_{l_{T}})\delta\big(t-(\tau_{l_{R}}+\tau_{l_{T}})\big).$
@@
Second, recall that the model in (6) is an expanded representation of the general sensing model in (4). The connection between the single path index $l$ in (4) and the index pair $(l_{T},l_{R})$ in (6) is given by the mapping $l=L^{\prime}(l_{T}-1)+l_{R}$ Therefore, by combining these steps, the definitive relationship between the parameters of the $l$th sensing path and its constituent communication paths can be written as
[…]
Note that this extrapolation is universally applicable since the underlying relationships in (30) are independent of the target’s physical properties. As shown in (28), the angle-dependent reflection coefficient $\beta(\theta_{l_{T}},\theta_{l_{R}})$ models the target’s specific scattering characteristics (e.g., shape or RCS) and acts as a component of the total complex gain $\alpha_{l_{T},l_{R}}$. However, it does not change the fundamental geometric parameters (angles $\theta_{l_{T}},\theta_{l_{R}}$ and delays $\tau_{l_{T}},\tau_{l_{R}}$), which are dictated solely by the propagation environment. Therefore, these relationships in (30) hold true for any passive sensing target at that location and enable extrapolation of sensing likelihood function $P(\hat{\mathbf{z}}|\mathbf{x})$ from the communication channel knowledge distribution $P(\mathbf{w}|\mathbf{q})$, regardless of its specific RCS or scattering characteristics.
@@ IV-B CADM-enabled sensing
$=$ $\sum_{l^{\prime}=1}^{L^{\prime}}\ln\big(\mathcal{N}(\theta_{l^{\prime}};\mu_{\theta,{l^{\prime}}}(\mathbf{q}),\sigma_{\theta,{l^{\prime}}}^{2}(\mathbf{q}))\mathcal{N}(\tau_{l^{\prime}};\mu_{\tau,{l^{\prime}}}(\mathbf{q}),\sigma_{\tau,{l^{\prime}}}^{2}(\mathbf{q}))\big)$
where $\mu_{\theta,l^{\prime}}(\mathbf{q})$ and $\sigma_{\theta,l^{\prime}}^{2}(\mathbf{q})$ are the mean and variance of the AoA $\theta_{l^{\prime}}$, and $\mu_{\tau,l^{\prime}}(\mathbf{q})$ and $\sigma_{\tau,l^{\prime}}^{2}(\mathbf{q})$ are the mean and variance of the delay $\tau_{l^{\prime}}$ for the $l^{\prime}$th path at UE location $\mathbf{q}$. These parameters are location-dependent, reflecting the unique channel characteristics at each $\mathbf{q}$ in the environment. The number of paths $L^{\prime}$ is chosen to capture the dominant propagation paths, with more complex environments requiring a larger $L^{\prime}$. The location-specific probabilistic modeling in (36) is crucial for the framework’s effectiveness in dynamic environments. While dominant propagation paths are determined by large and fixed structures, the movement of non-target objects (e.g., other vehicles or pedestrians) introduces statistical fluctuations. The learned variance captures this variations from historical data; a path that is frequently perturbed by moving scatterers will exhibit a larger spread in the training data, for which the model learns a higher variance. In this way, the typical environment dynamics are already embedded within the probabilistic CKM. The framework thus assumes a statistically stable environment, rather than a perfectly static one. This implies an environment where large-scale features (e.g., buildings, roads) are stable, while the variance learned by CKM accounts for the statistical fluctuations caused by dynamic elements such as vehicles and pedestrians. This assumption aligns well with typical urban environments.
@@
$=P_{\mathcal{M}}(\mathbf{w}_{l^{\prime}}=(\theta_{l_{T}},\tau_{l_{T}})|\mathbf{q}=\mathbf{x}),$
$=\mathcal{N}(\theta_{l_{T}};\mu_{\theta,{l_{T}}}(\mathbf{x}),\sigma_{\theta,{l_{T}}}^{2}(\mathbf{x}))\mathcal{N}(\tau_{l_{T}};\mu_{\tau,{l_{T}}}(\mathbf{x}),\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x}))$
@@
$=\mathcal{N}(\theta_{l_{R}};\mu_{\theta,{l_{R}}}(\mathbf{x}),\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x}))\mathcal{N}(\tau_{l_{R}};\mu_{\tau,{l_{R}}}(\mathbf{x}),\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x})).$
Then, the log-likelihood of the sensing channel knowledge can be written as (40) shown on the top of the next page, where the third equation is due to (32). In (38) and (39), it assumes that for a given path $l^{\prime}$, the angle $\theta_{l^{\prime}}$ and delay $\tau_{l^{\prime}}$ are statistically independent. This assumption is based on the underlying physics of multipath propagation in complex environments. From a physical standpoint, the parameters are determined by different geometric properties: The AoA is primarily dictated by the geometric orientation of the final scatterer relative to the receiver array. The path delay, in contrast, is a function of the entire path length traveled by the signal, which may involve multiple different scattering objects before reaching that final scatterer. In the NLoS multipath environment, the factors contributing to the total path length (and thus delay) are numerous and distinct from the geometric angle of the final hop. Therefore, in complex environments, the statistical relationship between the total path length and the final arrival angle is generally weak. It is also noted that the CKM framework is not fundamentally restricted by this assumption. The model could be extended to a full multivariate Gaussian distribution.
@@
$=$ $\sum_{l=1}^{L}\ln\Big(\mathcal{N}\left(\phi_{l};\mu_{\theta,{l_{T}}}\left(\mathbf{x}\right),\sigma_{\theta,{l_{T}}}^{2}\left(\mathbf{x}\right)\right)\mathcal{N}\left(\theta_{l};\mu_{\theta,{l_{R}}}\left(\mathbf{x}\right),\sigma_{\theta,{l_{R}}}^{2}\left(\mathbf{x}\right)\right)\mathcal{N}\left(\tau_{l};\mu_{\tau,{l_{T}}}\left(\mathbf{x}\right)+\mu_{\tau,{l_{R}}}\left(\mathbf{x}\right),\sigma_{\tau,{l_{T}}}^{2}\left(\mathbf{x}\right)+\sigma_{\tau,{l_{R}}}^{2}\left(\mathbf{x}\right)\right)\Big).$
$\nabla_{\mathbf{x}}$ $(-\ln P(\hat{\mathbf{z}}|\mathbf{x}))=\nabla_{\mathbf{x}}\sum_{l=1}^{L}-\ln\left(\mathcal{N}(\phi_{l};\mu_{\theta,{l_{T}}}(\mathbf{x}),\sigma_{\theta,{l_{T}}}^{2}(\mathbf{x}))\mathcal{N}(\theta_{l};\mu_{\theta,{l_{R}}}(\mathbf{x}),\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x}))\mathcal{N}(\tau_{l};\mu_{\tau,{l_{T}}}(\mathbf{x})+\mu_{\tau,{l_{R}}}(\mathbf{x}),\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})+\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x}))\right)$ (41)

### v3

@@ I Introduction
Significant progress has been made for ISAC systems, particularly in line-of-sight (LoS) scenarios, where the sensing targets are directly visible to both the transmitter and receiver. Various ISAC techniques have been developed, like joint waveform design, beamforming optimization, and interference management to balance radar-like sensing with data transmission, achieving high localization accuracy and spectral efficiency in LoS conditions [37, 29, 15, 46, 49]. For instance, methods such as time-of-arrival (ToA) estimation and angle-based localization leverage direct LoS paths to estimate target locations [36, 25]. However, in complex environments such as urban low-altitude airspace, LoS paths are frequently blocked. As a result, sensing in such environments can only rely on pure non-line-of-sight (NLoS) signal paths, posing a much greater challenge to map the estimated sensing parameters to the location of the sensing targets as more unknowns are involved [21, 26]. Some techniques have been proposed to achieve NLoS localization, including fingerprinting [31], single-base-station (BS) localization exploiting near-field scatterers [48], and reconfigurable intelligent surface (RIS)-aided methods [16]. However, they cannot be applied for sensing since unlike localization tasks, the sensing targets are typically non-cooperative in the sense that they only passively reflect the signals rather than transmitting or receiving them. Specifically, fingerprinting based localization works by building a database that maps user equipment (UE) location to its wireless communication channel signature such as received signal strength (RSS). However, this approach is unsuitable for sensing. The fundamental reason is that a fingerprint database characterizes the direct channel between UE and the anchor nodes, whereas sensing relies on the echoes from the passive sensing target. Therefore, the fingerprint database built for localization cannot be used for wireless sensing. Instead, sensing would require its own “sensing database” that fingerprints target-specific echo characteristics. Creating such databases is quite challenging, if not impossible, as different types of targets (e.g., vehicles, pedestrians) would need a distinct and comprehensive map of its potential echo signatures. On the other hand, single-BS near-field localization studied in [48] first senses near-field scatterers to acquire environment priors, and then leverages the location of near-field scatterers to localize UEs without relying on LoS link. However, it relies on near-field assumptions, limiting applicability in far-field ISAC. In addition, RIS-aided methods create virtual LoS relationships for localization by deploying reflective surfaces, they require additional costly and extensive hardware deployments. In fact, rigorous theoretical studies have proven that, without environment or NLoS prior knowledge, NLoS paths cannot provide beneficial information for localization [24]. This demonstrates the imperative need to develop an innovative framework that can leverage environment prior knowledge to tackle the challenging problem of NLoS sensing for ISAC systems.
@@
The remainder of this paper is organized as follows. Section II presents the system model for the ISAC framework. Section III discusses the motivation for CKM-enabled sensing, while Section IV details the mechanisms for the proposed CADM-enabled NLoS sensing, respectively. Section V provides simulation results, and Section VI concludes the paper.
@@ II System Model
where $c$ is the speed of light. As a result, after estimating the parameters $\hat{\theta}$ and $\hat{\tau}$ based on the echo signal $\mathbf{y}(t)$ in (1) with various algorithms reviewed in [3], the location of the sensing target $\mathbf{x}$ can be directly obtained as
@@
which ignores a significant number of potential paths. In contrast, the proposed model in (4) accounts for all $L=L^{\prime 2}$ possible path combinations, and the received signal in (4) can be written as
@@
To ultimately achieve the sensing goal, these parameters in $\mathbf{z}$ must be mapped to the location $\mathbf{x}$ of the sensing target. In LoS scenarios, this mapping can be directly achieved with the closed-form expression (3). However, such a mapping is no longer valid for the challenging scenario where LoS is absent, since NLoS paths dependent on the environment, denoted as $E$, as shown in Fig. 1. Therefore, when only NLoS paths are available, without any prior knowledge about the environment, determining the sensing target location $\mathbf{x}$ based on the sensed channel parameters $\mathbf{z}$ is an underdetermined problem. In fact, even in mixed scenarios with both LoS and NLoS paths, conventional methods typically employ techniques such as NLoS mitigation or path selection to isolate the LoS path, discarding NLoS paths that contain valuable target-related information [20]. This results in information loss, limiting the sensing accuracy. To utilize the target-related information in NLoS paths, which depends on $E$, an environment-aware sensing model that incorporates $E$ is necessary to leverage both LoS (if it exists) and NLoS paths for sensing target localization.
@@
The problem (10) can be solved by classic optimization methods such as gradient descent method. The update rule for the target location at the $(k+1)$th iteration is given by
@@ III Environment-aware ISAC via CKM
CKM was originally developed to infer prior channel knowledge based on UE locations for communication purposes. In this context, the sensing target location $\mathbf{x}$ was not involved. CKM offers a method to bypass the requirement for explicit modeling of $E$ or $f(\cdot)$, circumventing the complexities associated with environmental characterization and function derivation. Extending this framework to map sensing target locations $\mathbf{x}$ to channel knowledge $\mathbf{z}$ faces new challenges and opportunities, which motivates the exploration of CKM-enabled sensing in this work. In the following, we present the motivation behind this extension, detailing its core mechanism for enhancing environment-aware sensing capabilities.
Let $\mathbf{Q}=[\mathbf{q}_{1},\ldots,\mathbf{q}_{Q}]\in\mathbb{R}^{2\times Q}$ and $\mathbf{W}=[\mathbf{w}_{1},\ldots,\mathbf{w}_{Q}]\in\mathbb{R}^{D\times Q}$ denote the $Q$ historically visited UE locations and their corresponding observed $D$-dimensional channel knowledge data, respectively [42]. These location-channel knowledge pairs $(\mathbf{Q},\mathbf{W})$ are utilized to learn prior channel knowledge, forming the basis of CKM. CKM can thus be expressed as
@@
Figure 3: An illustration of CKM-enabled ISAC that utilizes the same CKM $\mathcal{M}$ developed for communication to achieve environment-aware sensing as well. The CKM, by learning the “location-to-channel” mapping ($P_{\mathcal{M}}(\mathbf{w}|\mathbf{q})$) from data, serves as an implicit representation of the environment $E$. This allows the framework to “skip” the computationally expensive step of explicit physical environment reconstruction and directly provide the sensing likelihood $P(\hat{\mathbf{z}}|\mathbf{x})$ for sensing target localization.
@@ IV-A Relationship between sensing response and communication channel
$\mathbf{y}_{l_{R},l_{T}}(t)$ $=\mathbf{h}_{R,l_{R}}(t)\ast\big(\beta(\theta_{l_{T}},\theta_{l_{R}})y_{T,l_{T}}(t)\big)$ (27)
$=\big(\alpha_{l_{R}}\mathbf{a}(\theta_{l_{R}})\delta(t-\tau_{l_{R}})\big)$
$\quad\ast\big(\beta(\theta_{l_{T}},\theta_{l_{R}})\alpha_{l_{T}}\mathbf{a}^{H}(\theta_{l_{T}})\mathbf{s}(t-\tau_{l_{T}})\big)$
$=\beta(\theta_{l_{T}},\theta_{l_{R}})\alpha_{l_{R}}\alpha_{l_{T}}\mathbf{a}(\theta_{l_{R}})$
$\quad\times\mathbf{a}^{H}(\theta_{l_{T}})\mathbf{s}\big(t-(\tau_{l_{R}}+\tau_{l_{T}})\big).$
@@
$=\sum_{l_{R}=1}^{L^{\prime}}\sum_{l_{T}=1}^{L^{\prime}}\beta(\theta_{l_{T}},\theta_{l_{R}})\alpha_{l_{R}}\alpha_{l_{T}}\mathbf{a}(\theta_{l_{R}})$
$\quad\times\mathbf{a}^{H}(\theta_{l_{T}})\mathbf{s}\big(t-(\tau_{l_{R}}+\tau_{l_{T}})\big)+\mathbf{n}(t)$
$=\mathbf{H}(t)\ast\mathbf{s}(t)+\mathbf{n}(t),$
@@
$\quad\times\mathbf{a}^{H}(\theta_{l_{T}})\delta\big(t-(\tau_{l_{R}}+\tau_{l_{T}})\big).$
@@
Second, recall that the model in (6) is an expanded representation of the general sensing model in (4). The connection between the single path index $l$ in (4) and the index pair $(l_{T},l_{R})$ in (6) is given by the mapping $l=L^{\prime}(l_{T}-1)+l_{R}$. Therefore, by combining these steps, the definitive relationship between the parameters of the $l$th sensing path and its constituent communication paths can be written as
$\phi_{l}$ $=\phi_{l_{T}}=\theta_{l_{T}},$ (30)
$\theta_{l}$ $=\theta_{l_{R}}=\theta_{l_{R}},$
$\tau_{l}$ $=\tau_{l_{T},l_{R}}=\tau_{l_{T}}+\tau_{l_{R}},$
where, for each row, the first, second, and third terms correspond to the parameters from (4), (6), and (29), respectively. Note that this extrapolation is universally applicable since the underlying relationships in (30) are independent of the target’s physical properties. As shown in (28), the angle-dependent reflection coefficient $\beta(\theta_{l_{T}},\theta_{l_{R}})$ models the target’s specific scattering characteristics (e.g., shape or RCS) and acts as a component of the total complex gain $\alpha_{l_{T},l_{R}}$. However, it does not change the fundamental geometric parameters (angles $\theta_{l_{T}},\theta_{l_{R}}$ and delays $\tau_{l_{T}},\tau_{l_{R}}$), which are dictated solely by the propagation environment. Therefore, these relationships in (30) hold true for any passive sensing target at that location and enable extrapolation of sensing likelihood function $P(\hat{\mathbf{z}}|\mathbf{x})$ from the communication channel knowledge distribution $P(\mathbf{w}|\mathbf{q})$, regardless of its specific RCS or scattering characteristics.
@@ IV-B CADM-enabled sensing
$=$ $\sum_{l^{\prime}=1}^{L^{\prime}}\ln\big(\mathcal{N}(\theta_{l^{\prime}};\mu_{\theta,{l^{\prime}}}(\mathbf{q}),\sigma_{\theta,{l^{\prime}}}^{2}(\mathbf{q}))$
$\times\mathcal{N}(\tau_{l^{\prime}};\mu_{\tau,{l^{\prime}}}(\mathbf{q}),\sigma_{\tau,{l^{\prime}}}^{2}(\mathbf{q}))\big)$
where $\mu_{\theta,l^{\prime}}(\mathbf{q})$ and $\sigma_{\theta,l^{\prime}}^{2}(\mathbf{q})$ are the mean and variance of the AoA $\theta_{l^{\prime}}$, and $\mu_{\tau,l^{\prime}}(\mathbf{q})$ and $\sigma_{\tau,l^{\prime}}^{2}(\mathbf{q})$ are the mean and variance of the delay $\tau_{l^{\prime}}$ for the $l^{\prime}$th path at UE location $\mathbf{q}$. These parameters are location-dependent, reflecting the unique channel characteristics at each $\mathbf{q}$ in the environment. The number of paths $L^{\prime}$ is chosen to capture the dominant propagation paths, with more complex environments requiring a larger $L^{\prime}$. The location-specific probabilistic modeling in (36) is crucial for the framework’s effectiveness in dynamic environments. While dominant propagation paths are determined by large and fixed structures, the movement of non-target objects (e.g., other vehicles or pedestrians) introduces statistical fluctuations. The learned variance captures these variations from historical data; a path that is frequently perturbed by moving scatterers will exhibit a larger spread in the training data, for which the model learns a higher variance. In this way, the typical environment dynamics are already embedded within the probabilistic CKM. The framework thus assumes a statistically stable environment, rather than a perfectly static one. This implies an environment where large-scale features (e.g., buildings, roads) are stable, while the variance learned by CKM accounts for the statistical fluctuations caused by dynamic elements such as vehicles and pedestrians. This assumption aligns well with typical urban environments.
@@
$=P_{\mathcal{M}}(\mathbf{w}_{l^{\prime}}=(\theta_{l_{T}},\tau_{l_{T}})|\mathbf{q}=\mathbf{x})$
$=\mathcal{N}(\theta_{l_{T}};\mu_{\theta,{l_{T}}}(\mathbf{x}),\sigma_{\theta,{l_{T}}}^{2}(\mathbf{x}))$
$\qquad\times\mathcal{N}(\tau_{l_{T}};\mu_{\tau,{l_{T}}}(\mathbf{x}),\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x}))$
@@
$=\mathcal{N}(\theta_{l_{R}};\mu_{\theta,{l_{R}}}(\mathbf{x}),\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x}))$
$\qquad\times\mathcal{N}(\tau_{l_{R}};\mu_{\tau,{l_{R}}}(\mathbf{x}),\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x})).$
Then, the log-likelihood of the sensing channel knowledge can be written as (40) shown on the top of the next page, where the second equation is due to (32). In (38) and (39), it assumes that for a given path $l^{\prime}$, the angle $\theta_{l^{\prime}}$ and delay $\tau_{l^{\prime}}$ are statistically independent. This assumption is based on the underlying physics of multipath propagation in complex environments. From a physical standpoint, the parameters are determined by different geometric properties: The AoA is primarily dictated by the geometric orientation of the final scatterer relative to the receiver array. The path delay, in contrast, is a function of the entire path length traveled by the signal, which may involve multiple different scattering objects before reaching that final scatterer. In the NLoS multipath environment, the factors contributing to the total path length (and thus delay) are numerous and distinct from the geometric angle of the final hop. Therefore, in complex environments, the statistical relationship between the total path length and the final arrival angle is generally weak. It is also noted that the CKM framework is not fundamentally restricted by this assumption. The model could be extended to a full multivariate Gaussian distribution.
@@
$=$ $\sum_{l=1}^{L}\ln\Big(\mathcal{N}\left(\phi_{l};\mu_{\theta,{l_{T}}}\left(\mathbf{x}\right),\sigma_{\theta,{l_{T}}}^{2}\left(\mathbf{x}\right)\right)\mathcal{N}\left(\theta_{l};\mu_{\theta,{l_{R}}}\left(\mathbf{x}\right),\sigma_{\theta,{l_{R}}}^{2}\left(\mathbf{x}\right)\right)$
$\times\mathcal{N}\left(\tau_{l};\mu_{\tau,{l_{T}}}\left(\mathbf{x}\right)+\mu_{\tau,{l_{R}}}\left(\mathbf{x}\right),\sigma_{\tau,{l_{T}}}^{2}\left(\mathbf{x}\right)+\sigma_{\tau,{l_{R}}}^{2}\left(\mathbf{x}\right)\right)\Big).$
$\nabla_{\mathbf{x}}(-\ln P(\hat{\mathbf{z}}|\mathbf{x}))=$ $\nabla_{\mathbf{x}}\sum_{l=1}^{L}-\ln\left(\mathcal{N}(\phi_{l};\mu_{\theta,{l_{T}}}(\mathbf{x}),\sigma_{\theta,{l_{T}}}^{2}(\mathbf{x}))\mathcal{N}(\theta_{l};\mu_{\theta,{l_{R}}}(\mathbf{x}),\sigma_{\theta,{l_{R}}}^{2}(\mathbf{x}))\right$. (41)
$\times\mathcal{N}(\tau_{l};\mu_{\tau,{l_{T}}}(\mathbf{x})+\mu_{\tau,{l_{R}}}(\mathbf{x}),\sigma_{\tau,{l_{T}}}^{2}(\mathbf{x})+\sigma_{\tau,{l_{R}}}^{2}(\mathbf{x}))\big)$